# Fundamental Oscillations of Massive Boson Stars - II

### v1

@@ 4.5 Detectability
We consider a model for burst GW as considered in [68], where oscillations are excited in a BS at $t=0$. We do not know any of any mechanisms that can excite $f$-mode oscillations for an isolated BS. The possible mechanisms to excite these could be i) during the formation of BS (like gravitational cooling) or ii) nearby astrophysical objects that hit the BS or interact closely (external perturbations). Here, we assume the oscillations are excited without delving deep into the excitation mechanism. We analyze the detectability assuming $f$-modes are excited. The oscillation frequency is its fundamental frequency $f$, which damps with a timescale of $\tau$ as it radiates GW. The GW strain amplitude $h(t)$ for such a signal is given by
@@
Here, $h_{0}$ is the peak amplitude measured by the GW detectors. It can be related to the total energy in the oscillations that is radiated away $E$, the distance of the detectors from the source of burst $d$, and the $f$-mode characteristics $f$ and $\tau$. This is given by [68, 69]
@@
The signal-to-noise ratio or SNR ($\rho$) for such a burst signal is derived as [70, 71]
@@
Here, $Q$ is the quality factor of the signal given by $Q=\pi f\tau$. This roughly gives the number of cycles ($N$) of the burst oscillations before the amplitude falls by a factor of $e$. We saw in Fig. 8 that $N_{min}$ for BSs is 406. This means that $Q\geq 406\pi$, which means $4Q^{2}\gg 1$. Thus, the SNR for BSs simply reduces to
@@
$S_{n}$ is the noise spectral density of the detector at the frequency of evaluation $f$.
We use Eq. 4.9 to calculate the SNR.
$\rho$ $=1.08\times 10^{-17}\left(\frac{1\mbox{ kpc}}{d}\right)\left(\frac{E}{M_{\odot}}\right)^{1/2}\left(\frac{1\mbox{ kHz}}{f}\right)\left(\frac{\mbox{Hz}^{-1/2}}{\sqrt{S_{n}}}\right)$ (4.10)
$=1.08\times 10^{-17}\left(\frac{1\mbox{ kpc}}{d}\right)\left(\frac{E}{M_{\odot}}\right)^{1/2}\frac{1}{202.67f^{\prime}(C)M^{\prime}(C)}\left(\frac{M_{BS}}{M_{\odot}}\right)\left(\frac{\mbox{Hz}^{-1/2}}{\sqrt{S_{n}}}\right)$ (4.11)
$\rho$ $=4.23\times 10^{-23}E_{d}\left(\frac{M_{BS}}{M_{\odot}}\right)^{3/2}\left(\frac{0.21}{f^{\prime}(C)}\right)\left(\frac{0.06}{M^{\prime}(C)}\right)\left(\frac{\mbox{Hz}^{-1/2}}{\sqrt{S_{n}}}\right)~{}.$ (4.12)
We have used the relation reported in [34] to write $f$ in terms of $M$ and $C$. This relation of BS SNR is completely in terms of $E$, $d$, $M$, and $C$ and does not include any explicit dependence on the $f$-mode characteristics. We have defined the parameter $E_{d}$ as
$E_{d}\equiv\frac{\sqrt{E/(10^{-10}M_{BS}})}{d/(1\text{kpc})}~{}.$ (4.13)
We have used a factor of $10^{-10}M_{BS}$ as a normalization of the energy emitted by the $f$-modes. Thus, $E_{d}=1$ means an energy corresponding to $10^{-10}$ times that of the mass-energy of the BS is radiated at a distance of $1$ kpc. Or alternatively, an energy or of $n\times 10^{-10}M_{BS}$ radiated at a distance of $\sqrt{n}$ kpc. We also note that the SNR for a fixed $E_{d}$ grows with $M_{BS}$. The primary reason for this is that the $f$-mode frequency for fixed compactness goes as the inverse of mass, and lower frequency results in larger signal strength.
In a realistic scenario, BSs in the universe might have different compactness depending on the initial formation conditions, but the model parameters are expected to be fixed. Thus, DM would have a fixed value of $x$, and compactness could be anything depending on the mass of the BS that formed. Hence, it is more natural to think in terms of $C$-$x$ parameters. Hence, we also provide the expression for $\rho(x,C)$ here. We can write $M_{BS}=M^{\prime}(C)xM_{Pl}^{3}$. Doing this we get
@@
The equality gives the minimum values of $E_{d}$ ($(E_{d})_{min}$) required for the BS burst source of compactness $C$ made of DM with parameter $x$ that is detectable with an SNR of 5 by a GW detector with the noise spectral density $S_{n}$. The lower the value of $(E_{d})_{min}$, the better the detectability. We expect this to be better for LISA as the SNR is higher for lower frequency (or higher mass).
@@
In Fig. 12(a) we plot this $(E_{d})_{min}$ (shown in colour) as a function of $x$ and $C$ for the case of advanced LIGO (aLIGO). $x$ and $(E_{d})_{min}$ are shown on log scales. The SNR and $(E_{d})_{min}$ depends on $f$. $f$ in turn depends on both $x$ and $C$. Since we are plotting $x$ on a log scale, the frequency, too, varies on a log scale here along the x-axis and increases along decreasing $x$ and increasing $C$. The parameter space shown for aLIGO covers frequencies in the range of 5-5000 Hz. We observe that detectability is better for lower frequencies (high $x$ and low $C$). The dark strip seen at the bottom right is due to the higher noise for aLIGO around 9.7 Hz. This worsens the detectability. We observe very high values for $(E_{d})_{min}$. $\log_{10}(E_{d})_{min}\gtrsim 9$ in the best case scenario.
@@
Figure 12: The detectability of BSs using (a) aLIGO and (b) LISA GW detectors. The colours indicate the minimum values of $E_{d}$ required for the detection with $\rho\geq 5$. The sensitivity curve for aLIGO is obtained from https://dcc.ligo.org/LIGO-T1500293/public. For LISA, the sensitivity curve was constructed using the analytical fit provided in [72]. The range for $x$ probed is different as aLIGO noise spectral density used is in the range 5-5000 Hz and 0.1 mHz - 1 Hz for LISA.
As the SNR is better for lower frequencies and higher BS mass, the detectability is expected to be the best in the case of LISA, as it is planned to be observed in the millihertz range. We make the same plot for the LISA detector (see Fig. 12(b)). The range of $x$ is such the frequencies fall in the LISA band, i.e., 0.1 mHz - 1Hz. We see that in this range $\log_{10}(E_{d})_{min}\gtrsim 4$. This means that for $E=10^{-10}M_{\odot}$, the distance of BS has to be $x\lesssim 0.1$ pc, i.e., roughly the size of Oort cloud. In the theoretical limiting case of $E=M_{BS}$, we reach the approximate distance of the galactic center from the Sun $\approx 10$ kpc. However, this is an unrealistic case. Thus, in the best possible case, we would be able to detect $f$-modes from BSs in the Solar System neighbourhood.
@@
The same range as that of aLIGO is probed by the CE and ET detectors. We expect better sensitivity with next-generation detectors as they have an order of magnitude improved sensitivity. We show the same plots of CE and ET in Fig. 13. A fainter band in the dark region on the CE plot is due to a dip in the noise at around 4kHz, improving the detectability. We also see a few very fine dark lines in the case of ET, all being due to narrow spikes in the noise spectral density. We see that detectabilities of CE and ET are comparable with $\log_{10}(E_{d})_{min}\gtrsim 7.8$. This improvement in order of magnitude compared to aLIGO is expected as $(E_{d})_{min}\propto\sqrt{S_{n}}$, and $\sqrt{S_{n}}$ is lower by an order of magnitude. Thus, in the LIGO frequency range corresponding to stellar mass BS, we have $\log_{10}(E_{d})_{min}\gtrsim 7.8$. Overall, these are very large values for $(E_{d})_{min}$ from a detectability point of view. $(E_{d})_{min}=8$ means for $E=10^{-10}M_{\odot}$, the distance to the BS has to be less than $10^{-8}$ kpc or $\approx 2$AU. This means detectability only extends to near-Earth BS or to BSs within the solar system for higher burst energies. For these detectors, we have seen that relevant BS masses are 0.1-10 $M_{\odot}$. In the theoretical limiting case of $E=M_{BS}$, which is unrealistic, we reach the distance of the nearest star $\approx 1$pc.
@@
Figure 13: The detectability of BSs using (a) Cosmic Explorer and (b) Einstein Telescope GW detectors. The colours indicate the minimum values of $E_{d}$ required for the detection with $\rho\geq 5$. The sensitivity curve for CE is obtained from https://dcc.cosmicexplorer.org/CE-T2000017/public and from https://dcc.ligo.org/LIGO-T1500293/public for ET. The range for $x$ probed is different as CE noise spectral density used is in the range 5-5000 Hz and 1 Hz - 10 kHz for ET.
@@ 5.1 Summary
In this work, we have focussed on a class of dark stars called dark stars made of scalar bosons. This is modelled by a scalar field $\phi$ that interacts gravitationally with all matter and with itself depending on the potential $V(\phi)$. We focus on the simplest potential given by $V(\phi)=\frac{1}{2}m^{2}\phi^{2}+\frac{1}{4}\lambda\phi^{4}$. The strong-interaction limit ($\Lambda=\lambda M_{Pl}^{2}/4\pi m^{2}/\gg 1$) results in the formation of a “massive star" whose mass scales as $\propto\sqrt{\lambda}M_{Pl}^{3}/m^{2}$ resembling massive fermionic stars. We study the non-radial $f$-mode oscillations of such massive BSs and the implications of scaling relations reported in [34].
@@
Scalar DM parameter space: We display the entire parameter space available for massive BSs and review all the available and relevant constraints. The astrophysical observations provide upper limits on $\lambda$ as a function of mass. A recent model independent study [40] gave a lower limit of scalar DM mass. Moreover, We stay in the $\Lambda\gg 1$ (put lower bound on $\lambda$ for the scalar DM to form massive BSs.
@@
Static properties: Mass and radius of massive BSs scale with the parameter $x=\sqrt{\lambda}/m^{2}$. Thus, relations between scaled mass $M^{\prime}$, scaled radius $R^{\prime}$, and $C^{\prime}=C$ are independent of the model parameters. We provide analytical fit functions (quadratic and quartic polynomials) along with fit coefficients and errors for the $M^{\prime}-C^{\prime}$, $R^{\prime}-C^{\prime}$, and $M^{\prime}-R^{\prime}$ curves and report the corresponding errors. The quartic fits are more accurate as expected, with an accuracy of $\lesssim 0.75\%$. The fits are very useful for future works of massive BS as direct analytical expressions can be used to calculate static observables instead of numerical solutions. We also find that the maximum density that can be reached at the centre of a stable BS is $\rho^{\prime}_{c}=1.6$.
@@
$f$-modes: We then evaluate the $f$-mode characteristics for massive BSs using a full general-relativistic setup. We use the $f$-mode frequency and damping time scaling relations as reported in [34]. We report analytical fits for the $f^{\prime}-M^{\prime}$, $f^{\prime}-C^{\prime}$, and $\tau^{\prime}-M^{\prime}$, $\tau^{\prime}-C^{\prime}$ relations, which can be used to evaluate the $f$-mode characteristics for any parameter values in the available parameter space along with the corresponding errors. All the fits are accurate to within $5\%$ and can be directly used in future works concerning $f$-modes of massive BSs without the need for any numerical calculations.
@@
Empirical relations: We then study the known empirical and quasi-universal relations for $f$-modes that are essential to probe the properties of interior matter. We showed that since scaling takes away the EoS dependency, we get a unique relation for $f^{\prime}=f^{\prime}(\sqrt{M^{\prime}/R^{\prime 3}})$ and $R^{\prime 4}\tau^{\prime}/M^{\prime 3}=\tau^{\prime}(M^{\prime}/R^{\prime})$. We provide linear fits to both that agree within $3\%$. The $\tau^{\prime}$ relation is found to be universal for BSs; however, it deviates from that for NSs.
@@
Universal relations: We also check the well-known universal relation of the mass-scaled $f$-mode characteristics with the dimensionless tidal deformability and compactness. These relations are well-studied for NSs [58, 59, 60, 73, 74], as well as NSs containing hyperons [75, 61, 76], deconfined quark matter [77, 78, 79, 80, 81], and DM-admixed NSs [82, 83, 84, 85, 86]. All these quantities, $M^{\prime}f^{\prime}$, $M^{\prime}/\tau^{\prime}$, $\Lambda$, $C$ do not scale with $x$. Hence, we get unique $M^{\prime}\omega_{f}-\Lambda$ and $M^{\prime}\omega_{f}-C$ for both the real and imaginary parts of the complex eigenfrequency, making it an exact relation for BSs. We report these relations here, along with corresponding analytical fits and errors. We also compare it with NS fits and find that the fits involving compactness deviate for BSs. We also report the fit for $N=f\tau=f^{\prime}\tau^{\prime}$ as a function of compactness, representing the number of cycles before the QNM damps $e$-folds.
@@
Observable parameter space: Focussing on the observational aspects of these $f$-modes for BSs shows the maximum possible $f$-mode frequency for each set of parameters $(\lambda,m)$ of DM, spanning the entire available parameter space. Using the frequency range of current and future GW detectors, we show what DM parameters space can be probed by the LISA, LIGO, and NEMO detectors. The space probed by ET and CE would span the region probed by LIGO and NEMO. We derived an upper limit of $x$ given by $x_{thresh}=2.66\times 10^{2}$ MeV2 such that model parameters with $x>x_{thresh}$ cannot be probed by any GW detector.
@@
Detectability: To check the detectability, we explore the $C-x$ parameters space, which represents a BS of compactness $C$ given the DM model parameters. For each point in this space, we calculate the lower bound on a quantity $E_{d}\equiv\sqrt{E}/d$ demanding that the SNR for the $f$-mode burst is $\gtrsim 5$. We find that for the $f$-mode energy of $E\approx 10^{-10}M_{BS}$, BSs within the solar system would be accessible to next-generation GW detectors like CE and ET. The detectability of BS $f$-modes is much better in the case of LISA, and BSs can be probed in the solar neighbourhood.

### v3

@@ 4.5 Detectability
We now see the prospects of detectability of these $f$-modes from massive BSs. Since these are quasinormal oscillations which damp out due to GW emission, we assume a model for burst GW as considered in [Ho2020], regardless of the damping time, where oscillations are excited in a BS at $t=0$. The possible mechanisms to excite these could be i) during the formation of a BS (like gravitational cooling) or ii) nearby astrophysical objects that hit the BS or interact closely (external perturbations). We analyze the detectability assuming $f$-modes are excited. The oscillation frequency is its fundamental frequency $f$, which damps with a timescale of $\tau$ as it radiates GWs. The GW strain amplitude $h(t)$ for such a signal is given by
@@
Here, $h_{0}$ is the peak amplitude measured by the GW detectors. It can be related to the total energy in the oscillations that is radiated away $E$, the distance of the detectors from the source of burst $d$, and the $f$-mode characteristics $f$ and $\tau$. This is given by [Ho2020, PradhanPathak2023]
@@
The signal-to-noise ratio or SNR ($\rho$) for such a burst signal is derived as [Echeverria1989, Kokkotas2001]
@@
Here, $Q$ is the quality factor of the signal given by $Q=\pi f\tau$. This roughly gives the number of cycles ($N$) of the burst oscillations before the amplitude falls by a factor of $e$. We saw in Fig. 5(c) that $N_{\rm min}$ for BSs is 406. This means that $Q\geq 406\pi$, which means $4Q^{2}\gg 1$. Thus, the SNR for BSs simply reduces to
@@
$S_{n}$ is the noise spectral density of the detector at the frequency of evaluation $f$. We use Eq. 4.9 to calculate the SNR.
$\rho$ $=4.23\times 10^{-23}E_{d}\left(\frac{M_{\rm BS}}{M_{\odot}}\right)^{3/2}\left(\frac{0.21}{f^{\prime}(C)}\right)\left(\frac{0.06}{M^{\prime}(C)}\right)\left(\frac{\mbox{Hz}^{-1/2}}{\sqrt{S_{n}}}\right)~.$ (4.10)
We have used the relation reported in [shirke_BS_fmodes_short] to write $f$ in terms of $M$ and $C$. This relation of BS SNR is completely in terms of $E$, $d$, $M$, and $C$ and does not include any explicit dependence on the $f$-mode characteristics. We have defined the parameter $E_{d}$ as
$E_{d}\equiv\frac{\sqrt{E/(10^{-10}M_{\rm BS}})}{d/(1\text{kpc})}~.$ (4.11)
We have used a factor of $10^{-10}M_{\rm BS}$ as a normalization of the energy emitted by the $f$-modes. Thus, $E_{d}=1$ means an energy corresponding to $10^{-10}$ times that of the mass-energy of the BS is radiated at a distance of $1$ kpc. Or alternatively, an energy or of $n\times 10^{-10}M_{\rm BS}$ radiated at a distance of $\sqrt{n}$ kpc. We also note that the SNR for a fixed $E_{d}$ grows with $M_{\rm BS}$. The primary reason for this is that the $f$-mode frequency for fixed compactness goes as the inverse of mass, and lower frequency results in larger signal strength.
In a realistic scenario, BSs in the universe might have different compactness depending on the initial formation conditions, but the DM parameters are expected to be fixed. Thus, DM would have a fixed value of $x$, and compactness could be anything depending on the mass of the BS that formed. Hence, it is more natural to think in terms of $C$-$x$ parameters. So we also provide the expression for $\rho(x,C)$ here. We can write $M_{\rm BS}=M^{\prime}(C)xM_{\rm Pl}^{3}$. Doing this we get
@@
The equality gives the minimum values of $E_{d}$ ($(E_{d})_{\rm min}$) required for the BS burst source of compactness $C$ made of DM with parameter $x$ that is detectable with an SNR of 5 by a GW detector with the noise spectral density $S_{n}$. The lower the value of $(E_{d})_{\rm min}$, the better the detectability.
@@
In Fig. 8(a) we plot this $(E_{d})_{\rm min}$ as a function of $x$ and $C$ for the case of advanced LIGO (aLIGO). We observe that detectability is better for lower frequencies (high $x$ and low $C$). The dark strip seen at the bottom right is due to the higher noise for aLIGO around 9.7 Hz. This worsens the detectability. We observe that $\log_{10}(E_{d})_{\rm min}>-3$ in the best case scenario. Thus, for $E=10^{-10}M_{\rm BS}$, BSs up to about 1 Mpc would be observable. Thus, aLIGO would be sensitive to BSs within and around our galaxy. For sources with $(x,C)$ such that the $f$-mode frequency is at the extreme limits of the observing range, BSs only within 100 pc ($\log_{10}(E_{d})_{\rm min}\sim 1$) would be observable.
@@
Figure 8: The detectability of BSs using (a) aLIGO and (b) LISA GW detectors. The colours indicate the minimum values of $E_{d}$ required for the detection with $\rho\geq 5$. The sensitivity curve for aLIGO is obtained from https://dcc.ligo.org/LIGO-T1500293/public. For LISA, the sensitivity curve was constructed using the analytical fit provided in [Robson2019]. The range for $x$ probed is different, as the noise spectral density range used is 5-5000 Hz for aLIGO and 0.1 mHz - 1 Hz for LISA.
As the SNR is better for lower frequencies and higher BS mass, the detectability is expected to be the best in the case of LISA, as it is planned to be observed in the millihertz range. We make the same plot for the LISA detector (see Fig. 8(b)). We see that LISA will be able to probe about a hundred times deeper with $\log_{10}(E_{d})_{\rm min}\gtrsim-5.5$. This means that for $E=10^{-10}M_{\rm BS}$, the distance of the BS has to be within $\sim 300$ Mpc. Since these are cosmological distances associated with burst energies, this corresponds to the luminosity distance. The BS masses probed in this case are over $100M_{\odot}$. For the extreme case of $M_{\rm BS}\approx 100M_{\odot}$, corresponding to $x=10^{-2}$MeV-2, the $f$-mode frequency lies at the higher frequency end of 1 Hz of LISA, and the observable distance is only about a kpc ($\log_{10}(E_{d})_{\rm min}\sim 0$).
@@
The same frequency range as that of aLIGO is probed by the CE and ET detectors. We expect better sensitivity with next-generation detectors as they have an order of magnitude improved sensitivity. We show the same plots of CE and ET in Fig. 9. A fainter band in the dark region on the CE plot is due to a dip in the noise at around 4kHz, improving the detectability. We also see a few very fine dark lines in the case of ET, all being due to narrow spikes in the noise spectral density. We see that detectabilities of CE and ET are comparable with $\log_{10}(E_{d})_{\rm min}\gtrsim-4.5$, which corresponds to a distance of $\sim 30$ Mpc for $E=10^{-10}M_{\rm BS}$. This improvement in more than an order of magnitude compared to aLIGO is expected as i) $(E_{d})_{\rm min}\propto\sqrt{S_{n}}$, and $\sqrt{S_{n}}$ is lower by an order of magnitude for ET and CE for $f\gtrsim 100$ Hz and ii) the SNR is higher for lower frequencies, and the sensitivity at low frequencies $f\lesssim 50$ Hz is better by more than an order of magnitude for CE and ET. For these detectors, we have seen that relevant BS masses are 0.1-10 $M_{\odot}$. For BSs with masses (and frequencies) on extreme ends of the frequency bands, the detectability is again about $\log_{10}(E_{d})_{\rm min}\sim 1$ corresponding to 100 pc. All the distances quoted are only for the assumed energy. For bursts with higher energies, farther distances can be probed and vice versa.
@@
Figure 9: The detectability of BSs using (a) Cosmic Explorer and (b) Einstein Telescope GW detectors. The colours indicate the minimum values of $E_{d}$ required for the detection with $\rho\geq 5$. The sensitivity curve for CE is obtained from https://dcc.cosmicexplorer.org/CE-T2000017/public and from https://dcc.ligo.org/LIGO-T1500293/public for ET. The range for $x$ probed is different as the CE noise spectral density used is in the range 5-5000 Hz and 1 Hz - 10 kHz for ET.
@@ 5.1 Summary
In a recent work [shirke_BS_fmodes_short], we investigated $f$-modes of massive BSs considering the $\phi^{4}$ self-interaction potential and demonstrated that they obey scaling relations independent of DM model parameters. In this work, we apply these relations to obtain useful analytical fits for global properties and universal relations important for GW asteroseismology, and comment on how to infer microscopic DM parameters with future GW observations. These results are summarized below.
@@
Scalar DM parameter space: We display the entire parameter space available for massive BSs and review all the available and relevant constraints. The astrophysical observations provide upper limits on $\lambda$ as a function of mass. A recent model-independent study [Zimmerman2024] gave a lower limit of the scalar DM mass. As we focus on massive BSs in the strong-interaction limit in this study, we stay in the $\Lambda>1000$ regime, which defines a lower bound on $\lambda$ for a given scalar DM mass.
@@
Static properties: Mass and radius of massive BSs scale with the parameter $x=\sqrt{\lambda}/m^{2}$. We provide analytical fit functions (quadratic and quartic polynomials) along with fit coefficients and errors for the known $M^{\prime}-C^{\prime}$, $R^{\prime}-C^{\prime}$, and $M^{\prime}-R^{\prime}$ curves and report the corresponding errors. The quartic fits have an accuracy of $\lesssim 0.75\%$. The fits are very useful for future works on massive BSs, as direct analytical expressions can be used to calculate static observables instead of numerical solutions, saving computation time. We also find that the maximum density that can be reached at the centre of a stable BS is $\rho_{c}=\rho^{\prime}_{c}/x^{2}=1.6/x^{2}$.
@@
$f$-modes: We then evaluate the $f$-mode characteristics for massive BSs using a full general-relativistic setup. We use the $f$-mode frequency and damping time scaling relations as reported in [shirke_BS_fmodes_short]. We report analytical fits for the $f^{\prime}-M^{\prime}$, $f^{\prime}-C^{\prime}$, and $\tau^{\prime}-M^{\prime}$, $\tau^{\prime}-C^{\prime}$ numerical relations, which can be used to evaluate the $f$-mode characteristics for any parameter values in the available parameter space. All the fits are accurate to within $5\%$ and can be directly used in future works concerning $f$-modes of massive BSs without the need for any numerical calculations.
@@
Empirical relations: We study the known empirical and quasi-universal relations for $f$-modes that are essential to probe the properties of interior matter. We showed that since scaling takes away the EoS dependency, we get a unique relation for $f^{\prime}=f^{\prime}(\sqrt{M^{\prime}/R^{\prime 3}})$ and $R^{\prime 4}\tau^{\prime}/M^{\prime 3}=\tau^{\prime}(M^{\prime}/R^{\prime})$. We provide linear fits to both that agree within $3\%$. The $\tau^{\prime}$ relation is found to be universal for BSs; however, it deviates from that for NSs.
@@
Universal relations: We also check the well-known universal relation of the mass-scaled $f$-mode characteristics with the dimensionless tidal deformability and compactness. We report unique relations $M^{\prime}\omega_{f}-\bar{\Lambda}$ and $M^{\prime}\omega_{f}-C$ for both the real and imaginary parts of the complex eigenfrequency, making it an exact relation for BSs, along with corresponding analytical fits and errors. We also compare it with NSs and find that the relations involving compactness deviate from those for BSs. We also report an analytical fit for $N=f\tau=f^{\prime}\tau^{\prime}$ numerical solution as a function of compactness, representing the number of cycles before the QNM damps $e$-folds.
@@
Observable DM parameter space: Focussing on the observational aspects of these $f$-modes for BSs, we show the maximum possible $f$-mode frequency for each set of parameters $(\lambda,m)$ of DM, spanning the entire available parameter space. Using the frequency range of current and future GW detectors, we show what DM parameter space can be probed by the LISA, LIGO, and NEMO detectors. We derive an upper limit of $x$ given by $x_{\rm thresh}=2.66\times 10^{2}$ MeV2 such that model parameters with $x>x_{\rm thresh}$ cannot be probed by any GW detector.
@@
Detectability: To check the detectability, we explore the CxC-x parameters space, which represents a BS of compactness $C$ given the DM model parameters. For each point in this space, we calculate the lower bound on a quantity $E_{d}\equiv\sqrt{E}/d$ assuming a GW burst model for all damping timescales, demanding that the SNR for the $f$-mode burst is $\gtrsim 5$. We find that for the $f$-mode energy of $E\approx 10^{-10}M_{\rm BS}$, GW detectors can probe BSs out to cosmological distances. In particular, aLIGO can detect $f$-modes of BSs up to a distance of 1 Mpc and up to 30 Mpc with the next-generation GW detectors like CE and ET. Higher distances, however, can still be probed for stronger bursts with larger $E$. The detectability of BS $f$-modes is much better in the case of LISA as compared to ET and CE, with an observable distance of about 300 Mpc.